OpenStudy (anonymous):

GIVING OUT A MEDAL What is the solution to g(x)= x2-2 step by step please :/

OpenStudy (anonymous):

sorry x^2-2

OpenStudy (johnweldon1993):

You want to find the x-intercepts I assume?

OpenStudy (anonymous):

idk it just says Find the solutions of g(x). Show each step.

OpenStudy (agent0smith):

Set it equal to zero first \[\Large x^2-2 = 0\]add 2 to both sides. Then take the square root of both sides.

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

\[X ^{2}-2=0\] add 2 too each side \[X ^{2}=2\] square root both sides \[X=\pm \sqrt{2}\] which is your answer \[X = \sqrt{2} (and) X= -\sqrt{2}\]